# Minimal model program for log canonical pairs on complex analytic spaces

### v1

@@ 4.2. Criterion for existence of log minimal model
Therefore, $N_{\sigma}(K_{\widetilde{X}}+\widetilde{\Delta})$ is well defined as an $\mathbb{R}$-divisor on $\widetilde{X}$, and this shows
@@
Thus $P_{\sigma}(K_{\widetilde{X}}+\widetilde{\Delta})$ is a finite $\mathbb{R}_{>0}$-linear combination of $\mathbb{Q}$-Cartier divisors on $\widetilde{X}$ that are nef over $W$. Moreover, by Lemma 3.4, we may replace $(X,\Delta)$ with $(\widetilde{X},\widetilde{\Delta})$. From these facts, by replacing $(X,\Delta)$ with $(\widetilde{X},\widetilde{\Delta})$, we may assume that
@@
such that $K_{\overline{X}}+\overline{\Delta}+\alpha\overline{M}$ is the limit of movable divisors over a neighborhood of $W$. Then
@@
for every prime divisor $\widetilde{D}$ on $\widetilde{X}$ whose image on $Y$ intersects $W$. By Lemma 3.4, the lc pair $(X,\Delta)$ (resp. $(X^{\prime},\Delta^{\prime})$) has a log minimal model over $Y$ around $W$ if and only if the lc pair $(\widetilde{X},\widetilde{\Delta})$ (resp. $(\widetilde{X},\widetilde{\Delta}^{\prime})$) has a log minimal model over $Y$ around $W$. Replacing $(X,\Delta)$ and $(X^{\prime},\Delta^{\prime})$ by $(\widetilde{X},\widetilde{\Delta})$ and $(\widetilde{X},\widetilde{\Delta}^{\prime})$ respectively, we may assume $X=X^{\prime}$ and that $(X,\Delta)$ and $(X^{\prime},\Delta^{\prime})$ are log smooth. Then [H20, Proof of Lemma 2.25] works with minor changes because we can use Theorem 4.16 instead of [H20, Theorem 2.23]. ∎
@@ 5.1. Auxiliary results
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,\Delta)$ be a dlt pair. Let $\mathcal{T}$ be a (possibly empty) finite set of prime divisors $D$ over $X$ such that the image of $D$ on $Y$ intersects $W$ and $-1<a(D,X,\Delta)<0$ for any $D\in\mathcal{T}$. For any open subset $U\subset Y$ containing $W$, we put $X_{U}=\pi^{-1}(U)$ and $\Delta_{U}=\Delta|_{X_{U}}$, and the restriction of any divisor $D\in\mathcal{T}$ to the inverse image of $U$ is denoted by $D_{U}$. Then there exist a Stein open subset $U\subset Y$ containing $W$, a projective bimeromorphism $f\colon\widetilde{X}\to X_{U}$ from a normal analytic variety $\widetilde{X}$, and $(\widetilde{X},\widetilde{\Delta})$ a disjoint union of dlt pairs satisfying the following.
@@
$-P_{\sigma}(f^{*}(K_{X}+\Delta);X^{\prime}/Y,W)\sim_{\mathbb{R},\,Z^{\prime}}-(F-N_{\sigma}(f^{*}(K_{X}+\Delta);X^{\prime}/Y,W))\geq 0.$
@@
      The pair $(\widetilde{X},\widetilde{\Delta}+t\widetilde{H})$ is dlt, $N_{\sigma}(K_{\widetilde{X}}+\widetilde{\Delta}+t\widetilde{H};\widetilde{X}/Y,W)$ is well defined as an $\mathbb{R}$-divisor on $\widetilde{X}$, and the support of $N_{\sigma}(K_{\widetilde{X}}+\widetilde{\Delta}+t\widetilde{H};\widetilde{X}/Y,W)$ does not depend on $t$, and
@@
      after shrinking $Y$ around $W$, the pair $(\widetilde{X},\widetilde{\Delta}-t\widetilde{G})$ has a good minimal model over $Y$ around $W$.
@@
    $X\dashrightarrow X_{1}$ is a sequence of steps of a $(K_{X}+\Delta+\lambda_{0}H)$-MMP over $Y$ around $W$ that is represented by a bimeromorphic contraction and $(X_{1},\Delta_{1}+\lambda_{0}H_{1})$ is a good minimal model of $(X,\Delta+\lambda_{0}H)$ over $Y$,
@@
    $X_{1}\dashrightarrow\cdots\dashrightarrow X_{i}\dashrightarrow\cdots$ is a sequence of steps of a $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$ with scaling of $\lambda_{0}H_{1}$ such that if we define
    $\lambda_{i}={\rm inf}\set{\mu\in\mathbb{R}_{\geq 0}\!}{\!K_{X_{i}}+\Delta_{i}+\mu H_{i}\text{\rm\;is nef over $W$}}$
    for each $i\geq 1$, then we have ${\rm lim}_{i\to\infty}\lambda_{i}=0$,
@@
over $U_{t}$ around $W$ to a good minimal model, which we may assume that the MMP is represented by a bimeromorphic contraction. By shrinking $Y$ to $U_{\lambda_{0}}$ and putting $X_{1}:=X^{\prime}_{\lambda_{0}}$, $\Delta_{1}:=\Delta^{\prime}_{\lambda_{0}}$, and $H_{1}:=H^{\prime}_{\lambda_{0}}$, we get a bimeromorphic contraction over $Y$
@@
By Proposition 5.5 and shrinking $Y$ around $W$, we get a dlt blow-up $(\widetilde{X},\widetilde{\Delta})\to(X,\Delta)$ and effective $\mathbb{R}$-Cartier divisors $\widetilde{G}$ and $\widetilde{H}$ on $\widetilde{X}$ such that $\widetilde{X}$ is $\mathbb{Q}$-factorial over $W$ and the properties (I), (II), (III), (III-a), and (III-b) in Proposition 5.5 hold. Then it is easy to check that we may replace $(X,\Delta)$ by $(\widetilde{X},\widetilde{\Delta})$. Therefore, replacing $(X,\Delta)$ by $(\widetilde{X},\widetilde{\Delta})$, we may assume that $X$ is $\mathbb{Q}$-factorial over $W$ and there exist effective $\mathbb{R}$-Cartier divisors $G$ and $H$ on $X$ satisfying (I), (II), (III), (III-a), and (III-b) in Proposition 5.5. By Proposition 5.6 and shrinking $Y$ around $W$, we obtain a real number $\lambda_{0}>0$, a bimeromorphic contraction
@@
    $X\dashrightarrow X_{1}$ is a sequence of steps of a $(K_{X}+\Delta+\lambda_{0}H)$-MMP over $Y$ around $W$ that is represented by a bimeromorphic contraction and $(X_{1},\Delta_{1}+\lambda_{0}H_{1})$ is a good minimal model of $(X,\Delta+\lambda_{0}H)$ over $Y$,
@@
    $X_{1}\dashrightarrow\cdots\dashrightarrow X_{i}\dashrightarrow\cdots$ is a sequence of steps of a $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$ with scaling of $\lambda_{0}H_{1}$ such that if we define
    $\lambda_{i}={\rm inf}\set{\mu\in\mathbb{R}_{\geq 0}\!}{\!K_{X_{i}}+\Delta_{i}+\mu H_{i}\text{\rm\;is nef over $W$}}$
    for each $i\geq 1$, then we have ${\rm lim}_{i\to\infty}\lambda_{i}=0$,
@@
From now on, we prove that for any integer $d\geq 0$, there is $m_{d}\geq m$ such that after $m_{d}$ steps, the non-biholomorphic locus of the $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$ does not intersect any $d$-dimensional lc center of $(X_{m_{d}},\Delta_{m_{d}})$ whose image on $Y$ intersects $W$. The case $d={\rm dim}\,X-1$ of this statement and (5) imply the termination of the $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$. Therefore, to prove Theorem 5.7 it is sufficient to prove the statement. We will prove the statement by induction on $d$. In the case where $d=0$, we can take $m$ as $m_{0}$. We assume the existence of $m_{d-1}$ in the statement. By replacing $m$ with $m_{d-1}$, we may assume $m_{d-1}=m$. By Theorem 3.17 and replacing $m$ again, we may assume that for any $d$-dimensional lc center $S_{m}$ of $(X_{m},\Delta_{m})$ whose image on $Y$ intersects $W$, the following condition holds: For every $i\geq m$, if we shrink $Y$ to $Y_{i}$ so that the $(K_{X_{1}}+\Delta_{1})$-MMP $(X_{1},\Delta_{1})\dashrightarrow(X_{i},\Delta_{i})$ is represented by a bimeromorphic contraction $X_{1}\dashrightarrow X_{i}$, then the induced bimeromorphic map $S_{m}\dashrightarrow S_{i}$ is small and the strict transform of $\Delta_{S_{m}}$ (resp. $H_{S_{m}}$) on $S_{i}$ is equal to $\Delta_{S_{i}}$ (resp. $H_{S_{i}}$). Since ${\rm lim}_{i\to\infty}\lambda_{i}=0$, by Theorem 3.18, to prove the existence of $m_{d}$ it is sufficient to prove that $(S_{m},\Delta_{S_{m}})$ has a log minimal model over $Y$ around $W$.
@@
In this step we apply [H20, Step 2 in the proof of Theorem 3.4] to our setting. Fix a $d$-dimensional lc center $S_{m}$ of $(X_{m},\Delta_{m})$ whose image on $Y$ intersects $W$. By replacing $Y$ with $Y_{m}$, we may assume that the $(K_{X_{1}}+\Delta_{1})$-MMP $(X_{1},\Delta_{1})\dashrightarrow(X_{m},\Delta_{m})$ is represented by a bimeromorphic contraction $X_{1}\dashrightarrow X_{m}$. Then $X\dashrightarrow X_{m}$ is also a bimeromorphic contraction. By construction of $X\dashrightarrow X_{m}$ in (1)–(3) in Step 1, we can find an lc center $S$ of $(X,\Delta)$ such that $X\dashrightarrow X_{m}$ induces a birmeromorphic map $S\dashrightarrow S_{m}$. Using the lc center $S$, in the rest of this step, we shrink $Y$ appropriately and we establish the following variety, divisor, and inequalities.
@@
    for every $i\geq m$, if we replace $Y$ by $Y_{i}$ so that the log MMP $(X_{1},\Delta_{1})\dashrightarrow(X_{i},\Delta_{i})$ is represented by a bimeromorphic contraction $X_{1}\dashrightarrow X_{i}$, then
    holds for any prime divisors $Q$ over $S$ whose image on $Y$ intersects $W$, and
@@
$\mathcal{C}=\Set{\bar{D}}{\begin{array}[]{l}\!\!\text{$\bar{D}$ is a prime divisor on $S$ such that the image of $\bar{D}$ on $Y$}\\ \!\!\text{intersects $W$ and $a(\bar{D},S_{m},\Delta_{S_{m}})<a(\bar{D},S,\Delta_{S})$}\end{array}\!\!}.$
@@
for all $\bar{D}\in\mathcal{C}$. Since every element of $\mathcal{C}$ is a prime divisor on $S$, we see that all elements of $\mathcal{C}$ are components of $H_{S}$. Thus $\mathcal{C}$ is a finite set, and furthermore, any $\bar{D}\in\mathcal{C}$ satisfies $-1<a(\bar{D},S,\Delta_{S}+\lambda_{m}H_{S})$. Thus, we obtain
$-1<a(\bar{D},S_{m},\Delta_{S_{m}})<0$
@@
[…]
$-1\leq a(Q^{\prime},S_{m},\Delta_{S_{m}})\leq a(Q^{\prime},T,\Psi)$
@@ 5.2. Proof of main results
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,\Delta)$ be an lc pair. Let $A$ be an effective $\mathbb{R}$-divisor on $X$ such that $(X,\Delta+A)$ is an lc pair and $K_{X}+\Delta+A$ is nef over $W$. Then no infinite sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of $A$
@@
  • If we define $\lambda_{i}={\rm inf}\set{\mu\in\mathbb{R}_{\geq 0}\!}{\!K_{X_{i}}+\Delta_{i}+\mu A_{i}\text{\rm\, is nef over }W}$, where $A_{i}$ is the strict transform of $A$ on $X_{i}$, then ${\rm lim}_{i\to\infty}\lambda_{i}=0$, and
  • there are infinitely many $i$ such that $(X_{i},\Delta_{i})$ is log abundant over $Y$.
@@
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,B+A)$ be an lc pair, where $B$ is an effective $\mathbb{R}$-divisor on $X$ and $A$ is an effective $\pi$-ample $\mathbb{R}$-divisor on $X$. Let $f\colon(\tilde{X},\tilde{B})\to(X,B)$ be a dlt blow-up of $(X,B)$, and we put $\tilde{\Gamma}=\tilde{B}+f^{*}A$. Let $\tilde{H}$ be a $(\pi\circ f)$-ample $\mathbb{R}$-divisor on $\tilde{X}$ such that $(\tilde{X},\tilde{\Gamma}+\tilde{H})$ is lc and $K_{\tilde{X}}+\tilde{\Gamma}+\tilde{H}$ is nef over $W$. Then there exists a sequence of steps of a $(K_{\tilde{X}}+\tilde{\Gamma})$-MMP over $Y$ around $W$ with scaling of $\tilde{H}$
@@
be a sequence of steps of a $(K_{\tilde{X}}+\tilde{\Gamma})$-MMP over $Y$ around $W$ with scaling of $\tilde{H}$. By the argument as in [H22a], we see that all $(\tilde{X}_{i},\tilde{\Gamma}_{i})$ are log abundant over $Y$. Moreover, if we define
$\lambda_{i}={\rm inf}\set{\mu\in\mathbb{R}_{\geq 0}\!}{\!K_{\tilde{X}_{i}}+\tilde{\Gamma}_{i}+\mu\tilde{H}_{i}\text{\rm\, is nef over }W},$
where $\tilde{H}_{i}$ is the strict transform of $\tilde{H}$ on $\tilde{H}_{i}$, then ${\rm lim}_{i\to\infty}\lambda_{i}=0$. By Theorem 5.8, the $(K_{\tilde{X}}+\tilde{\Gamma})$-MMP terminates. Therefore Theorem 5.9 holds. ∎

### v2

@@ 4.2. Criterion for existence of log minimal model
[…]
@@
[…]
@@
over $Y$ such that $K_{\overline{X}}+\overline{\Delta}+\alpha\overline{M}$ is the limit of movable divisors over a neighborhood of $W$. Then
@@
[…]
@@ 5.1. Auxiliary results
[…]
@@
$P_{\sigma}(f^{*}(K_{X}+\Delta);X^{\prime}/Y,W)\sim_{\mathbb{R},\,Z^{\prime}}-(F-N_{\sigma}(f^{*}(K_{X}+\Delta);X^{\prime}/Y,W))\geq 0.$
@@
[…]
@@
[…]
@@
    $X\dashrightarrow X_{1}$ is a sequence of steps of a $(K_{X}+\Delta+\lambda_{0}H)$-MMP over $Y$ around $W$ that is represented by a bimeromorphic contraction over $Y$ and $(X_{1},\Delta_{1}+\lambda_{0}H_{1})$ is a good minimal model of $(X,\Delta+\lambda_{0}H)$ over $Y$,
@@
    $X_{1}\dashrightarrow\cdots\dashrightarrow X_{i}\dashrightarrow\cdots$ is a sequence of steps of a $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$ with scaling of $\lambda_{0}H_{1}$ such that if we define λ_i=inf{μR_0 —  K_X_i+Δ_i+μH_i is nef over WW} for each $i\geq 1$, then we have ${\rm lim}_{i\to\infty}\lambda_{i}=0$,
@@
over $U_{t}$ around $W$ to a good minimal model, which we may assume that the MMP is represented by a bimeromorphic contraction over $U_{t}$. By shrinking $Y$ to $U_{\lambda_{0}}$ and putting $X_{1}:=X^{\prime}_{\lambda_{0}}$, $\Delta_{1}:=\Delta^{\prime}_{\lambda_{0}}$, and $H_{1}:=H^{\prime}_{\lambda_{0}}$, we get a bimeromorphic contraction over $Y$
@@
[…]
@@
    $X\dashrightarrow X_{1}$ is a sequence of steps of a $(K_{X}+\Delta+\lambda_{0}H)$-MMP over $Y$ around $W$ that is represented by a bimeromorphic contraction over $Y$ and $(X_{1},\Delta_{1}+\lambda_{0}H_{1})$ is a good minimal model of $(X,\Delta+\lambda_{0}H)$ over $Y$,
@@
    $X_{1}\dashrightarrow\cdots\dashrightarrow X_{i}\dashrightarrow\cdots$ is a sequence of steps of a $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$ with scaling of $\lambda_{0}H_{1}$ such that if we define λ_i=inf{μR_≥0 —  K_X_i+Δ_i+μH_i is nef over WW} for each $i\geq 1$, then we have ${\rm lim}_{i\to\infty}\lambda_{i}=0$,
@@
From now on, we prove that for any integer $d\geq 0$, there is $m_{d}\geq m$ such that after $m_{d}$ steps, the non-biholomorphic locus of the $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$ does not intersect any $d$-dimensional lc center of $(X_{m_{d}},\Delta_{m_{d}})$ whose image on $Y$ intersects $W$. The case $d={\rm dim}\,X-1$ of this statement and (5) imply the termination of the $(K_{X_{1}}+\Delta_{1})$-MMP over $Y$ around $W$. Therefore, to prove Theorem 5.7 it is sufficient to prove the statement. We will prove the statement by induction on $d$. In the case where $d=0$, we can take $m$ as $m_{0}$. We assume the existence of $m_{d-1}$ in the statement. By replacing $m$ with $m_{d-1}$, we may assume $m_{d-1}=m$. By Theorem 3.17 and replacing $m$ again, we may assume that for any $d$-dimensional lc center $S_{m}$ of $(X_{m},\Delta_{m})$ whose image on $Y$ intersects $W$, the following condition holds: For every $i\geq m$, if we shrink $Y$ to $Y_{i}$ so that the $(K_{X_{1}}+\Delta_{1})$-MMP $(X_{1},\Delta_{1})\dashrightarrow(X_{i},\Delta_{i})$ is represented by a bimeromorphic contraction $X_{1}\dashrightarrow X_{i}$ over $Y$, then the induced bimeromorphic map $S_{m}\dashrightarrow S_{i}$ is small and the strict transform of $\Delta_{S_{m}}$ (resp. $H_{S_{m}}$) on $S_{i}$ is equal to $\Delta_{S_{i}}$ (resp. $H_{S_{i}}$). Since ${\rm lim}_{i\to\infty}\lambda_{i}=0$, by Theorem 3.18, to prove the existence of $m_{d}$ it is sufficient to prove that $(S_{m},\Delta_{S_{m}})$ has a log minimal model over $Y$ around $W$.
@@
In this step we apply [H20, Step 2 in the proof of Theorem 3.4] to our setting. Fix a $d$-dimensional lc center $S_{m}$ of $(X_{m},\Delta_{m})$ whose image on $Y$ intersects $W$. By replacing $Y$ with $Y_{m}$, we may assume that the $(K_{X_{1}}+\Delta_{1})$-MMP $(X_{1},\Delta_{1})\dashrightarrow(X_{m},\Delta_{m})$ is represented by a bimeromorphic contraction $X_{1}\dashrightarrow X_{m}$ over $Y$. Then $X\dashrightarrow X_{m}$ is also a bimeromorphic contraction over $Y$. By construction of $X\dashrightarrow X_{m}$ in (1)–(3) in Step 1, we can find an lc center $S$ of $(X,\Delta)$ such that the map $X\dashrightarrow X_{m}$ induces a birmeromorphic map $S\dashrightarrow S_{m}$. Using the lc center $S$, in the rest of this step, we shrink $Y$ appropriately and we establish the following variety, divisor, and inequalities.
@@
    for every $i\geq m$, if we replace $Y$ by $Y_{i}$ so that the log MMP $(X_{1},\Delta_{1})\dashrightarrow(X_{i},\Delta_{i})$ is represented by a bimeromorphic contraction $X_{1}\dashrightarrow X_{i}$ over $Y$, then a(Q,S,Δ_S+λ_iH_S)≤a(Q,S_i,Δ_S_i+λ_iH_S_i) holds for any prime divisors $Q$ over $S$ whose image on $Y$ intersects $W$, and
@@
[…]
@@
for all $\bar{D}\in\mathcal{C}$. Since every element of $\mathcal{C}$ is a prime divisor on $S$, we see that all elements of $\mathcal{C}$ are components of $H_{S}$. Thus $\mathcal{C}$ is a finite set, and furthermore, any $\bar{D}\in\mathcal{C}$ satisfies $1<a(\bar{D},S,\Delta_{S}+\lambda_{m}H_{S})$. Thus, we obtain
$1<a(\bar{D},S_{m},\Delta_{S_{m}})<0$
@@
is $\mathbb{R}$-Cartier, where $D|_{T^{\prime}}=\bigcup_{j}Q_{D,\,j}$ is the prime decomposition. Then the images of some prime divisors $Q_{D,\,j}$ on $Y^{\prime}$ may not intersect $W$. However, by replacing $Y^{\prime}$ with a smaller Stein open subset to remove such $Q_{D,\,j}$, we may assume that $\Psi|_{T^{\prime}}$ is $\mathbb{R}$-Cartier. In particular, we may assume that $K_{T}+\Psi$ is $\mathbb{R}$-Cartier. Moreover we may assume that the image of any component of $K_{T}+\Psi-\psi^{*}(K_{S_{m}}+\Delta_{S_{m}})$ intersects $W$. By ($\star\star$2), we obtain $K_{T}+\Psi\leq\psi^{*}(K_{S_{m}}+\Delta_{S_{m}})$. From this, we have
$1\leq a(Q^{\prime},S_{m},\Delta_{S_{m}})\leq a(Q^{\prime},T,\Psi)$
@@ 5.2. Proof of main results
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,\Delta)$ be an lc pair. Let $A$ be an effective $\mathbb{R}$-divisor on $X$ such that $(X,\Delta+A)$ is an lc pair and $K_{X}+\Delta+A$ is nef over $W$. Then no infinite sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of AA
@@
[…]
@@
[…]
@@
[…]